# Hyperfine interaction in the Autler-Townes effect II: control of two-photon selection rules in the Morris-Shore basis

### 05 Dec 2023 v1

@@ I Introduction
Although a variety of coherent techniques such as the stimulated Raman adiabatic passage technique (STIRAP) Vitanov et al. (2017); Shore (2017) have enabled nearly $100\%$ efficient excitation Garcia-Fernandez et al. (2005), for alkali-metal atoms the perfect level addressing problem remains unsolved due to existence of Hyperfine (HF) splitting of the atoms’ energy levels. HF structure of highly excited levels with large principal quantum numbers $n$ can not be resolved due to the small energy separation ($\sim n^{-3}$) between sublevels. Thus no separate HF sublevel can be populated by any tuning of the exciting laser. Our research presented here (see Sections II and IV.2) demonstrates that selective excitation can still be achieved for a few two-photon excitation schemes of type $3(5)s_{1/2}-3(5)p_{3/2,1/2}-ns_{1/2}$ in ${}^{23}$a and ${}^{87}$b atoms (see Fig. 1). Throughout this paper, the principal quantum numbers and the HF energy level splitting values placed in parentheses refer to rubidium atoms. In these excitation schemes, the total angular momentum ${F}^{\prime\prime}$ of a well-resolved ground $3(5)s_{1/2}$ state has to remain the same for the final $ns_{1/2}$ state. As long as the P-laser couples all HF components $F^{\prime}$ of a $3(5)p_{3/2,1/2}$ state in the first excitation step ($\Delta F_{PS}=0,\pm 1$ according to the one-photon selection rule), while in the second step the S-laser couples both HF components $F=1,2$ of $ns_{1/2}$ state ($\Delta F_{SP}=0,\pm 1$), one of which turns out to be unpopulated, we can speak of a specific two-photon selection rule $\Delta F_{SPS}\equiv 0$ observed instead of the expected $\Delta F_{SPS}=0,\pm 1$.
@@ II HF blocks architecture of two-step excitation schemes
Dynamics of atomic systems under the influence of periodic or polychromatic control fields can be accurately described using various modifications of the Floquet technique Chu and Telnov (2004); Efimov et al. (2014); Eckardt and Anisimovas (2015). A useful tool for theoretical analysis in the case of nearly–resonant monochromatic laser fields $\textbf{E}\cos(\omega_{L}t)$ is the rotating wave approximation (RWA) Shore (2009). In this framework, the problems of light-matter interactions are reduced to solving the quasistationary optical Bloch equations for evolution of the atomic density matrix, with characteristic timescales being determined by the duration of laser pulses, i.e. the temporal behavior of the field amplitudes $\textbf{E}(t)$. A key component of the theoretical analysis is the concept of adiabatic (“laser-dressed”) atomic state basis, which often allows to develop quantitative description of the phenomena at hand for relatively simple systems and to predict possible evolution scenarios in the case of multidimensional configurations Shore (2017); Rangelov et al. (2006). The latter have been considered as a promising physical objects for quantum processing on the so-called qudits (multidimensional extension of qubits), permitting to design effective algorithms for the implementation of universal gates via adiabatic passage processes Rousseaux et al. (2013); Cinins et al. (2022). In this section we develop an approach that enables identifying dressed states in multidimensional two-step ladder excitation schemes, in the specific case of alkali metal atoms interacting with linearly polarized laser light.
@@ II.2 Dipole matrix elements
The relation (5) indicates that both laser fields couple only the levels with the same quantum numbers $m_{J},m_{I}$, and the corresponding linkage diagram shown in Fig. 3 is reduced to a set of separate, non-interacting three-level ladders
@@ II.3 Building the Morris-Shore basis vectors
The arbitrarily chosen basis vectors $|d_{M\xi}\rangle$ ($\xi\!=\!0,...,\kappa$) in the manifold $\Lambda_{nd_{5/2}}^{DS}$ are in no way coupled with the ground states (see Fig.2), i.e. are “Dark”.
@@ III.1 Formation of multiplet structure in AT signals
Figure 4: Energies $\varepsilon$ of adiabatic (dressed) states in Na vs the pump field Rabi frequency $\Omega_{S}$ for the $3s_{1/2}(F^{\prime\prime},M)\!\to\!3p_{3/2}(F^{\prime},M)\!\to\!4d_{5/2}(F,M)$ excitation Zeeman sequences ($M\!=\!0,1,2$) in the case of pump laser detuning $\Delta_{S}\!=\!-30\,\text{MHz}$. The dashed vertical lines correspond to $\Omega_{S}$-values presented in Fig. 5 frames (a)-(d). The square brackets indicate the states that transfer into dark states. The zero-energy positions related to the energy of the $3p_{3/2}(F^{\prime}\!=\!2)$-state. The curves labeled with quantum numbers $F^{\prime}$ or $F$ refer to the dressed states starting to evolve at $\Omega_{S}=0$ from the corresponding HF basis vectors ${\left|\gamma,F,M\right.\rangle}$ ($\gamma\!=\!i,f$).
@@
Strong coupling implies that both $\Omega_{1/2,3/2}^{(S)}$ exceed the bare energy separations $\Delta{\widetilde{\varepsilon}}_{\varsigma}$ (14) for all coupling pairs, so that at large $\Omega_{S}$ relation (14) reduces to the simple linear form:
@@
Figure 6: (Color online) The same as in Fig.4 in the case of $3s_{1/2}(F^{\prime\prime}=2)\to 3p_{3/2}\to 5s_{1/2}$ excitation sequence in Na for different $S$-field Rabi frequencies. The corresponding RWA linkage diagrams are shown in frames (a1), (b1), (c1). Simulations are performed using parameters $\Omega_{P}=0.87\,\text{MHz}$, $\Delta_{S}=-115\,\text{MHz}$.
@@ III.2 Numerical simulation
Figure 7: (Color online) The same as in Fig.5 in the case of $3s_{1/2}(F^{\prime\prime}=2)\to 3p_{3/2}\to 5s_{1/2}$ excitation sequence for different $S$ field Rabi frequencies. Simulations are performed for Na atoms using parameters $\Omega_{P}=0.87\,\text{MHz}$, $\Delta_{S}=-115\,\text{MHz}$. The expected AT peaks (the vertical lines) are consistent with the data in Fig.6.
@@ IV.1 Spectral composition and selection rules
With regard to the spectra of Figs. 5, 7, one cleanly sees both signs of spectra pauperisation, namely: (a) multiplet structure and (b) lack of all AT multiplets $1_{\pm}$ related to the excitation scheme $3s_{1/2}(F^{\prime\prime}\!=\!1)\rightarrow 3p_{3/2}$. Bright multiplets $2_{\pm}$ and $1_{\pm}$ turn out to complement each other, i.e. their visualization occurs during independent probing from different initial HF components $F^{\prime\prime}\!=\!2,1$ of the ground state. The data in Table 2(a) do show that HF interaction poorly couples all ${\left|i\right.\rangle}_{1}$ and ${\left|i\right.\rangle}_{2}$ BS: the corresponding $\Omega_{HF}$ rates are $\Omega_{HF}\!=\!0$ for $M\!=\!0,2$ while for two BS with $M\!=\!1$ the sharing parameter (17) $R\!=\!(7.44/65.6)^{2}\!=\!0.0129$. Practically the same thing happens in the case of Rb atoms, as can be seen from the corresponding data in Table 3(a): for both $M\!=\!0,2$ the values of $\Omega_{HF}\!=\!0$ and the the sharing parameter (17) $R\!=\!(34.0/299.7)^{2}\!=\!0.0128$.
@@ V Conclusions
An interesting feature of the excitation ladders $ks_{1/2}\to kp_{3/2,1/2}\to ns_{1/2}$ manifests as preservation of the HF quantum number $F$ in transitions between the initial $ks_{1/2}$ and the final $ns_{1/2}$ states, resulting in a specific selection rule $\Delta F\!=\!F^{\prime\prime}\!-\!F=0$ for the two-photon transitions $ks_{1/2}(F^{\prime\prime})\to ns_{1/2}(F)$. The reason for this phenomenon is related to the physics of the corresponding excitation processes proceeding through a series of partial two-photon paths with different intermediate HF sublevels (see Fig. 6 (b1)). The destructive interference between probability amplitudes forbids the two-photon transition when $F_{f}\neq F_{g}$. This observation opens practically important perspectives for artificially introducing custom selection rules in many-photon interactions both to manipulate quantum states and to achieve individual addressing of HF components of atomic and molecular energy levels.
@@
In the absence of HF interaction, the control radiation of P- and S-lasers allows ideal ($100\%$) selective excitation of single three-level ladder blocks (10), i.e. excitation of selected bright $\eta$-states with a fixed $\eta$-index ($\eta=F^{\prime\prime}M$). With them, for example, using the STIRAP technique, it is possible to carry out some basic quantum operations Vitanov et al. (2017) or, on the basis of these three $\eta$-states, to create independent cells for storing optical information by forming proper $\eta$-polaritons Vitanov et al. (2017). Coherent excitation of Rb Rydberg states in a three-level Levine et al. (2019) or four-level Beterov et al. (2023) elementary ladder blocks with F”=2, M=0 is an important element of a three-qubit Toffoli gate implementation Levine et al. (2019). The data presented in Table 3 provides useful information for determining the optimal parameters (Rabi frequencies and laser detunings) in this type of experiments Levine et al. (2019); Beterov et al. (2023) which use high-contrast Rydberg pulses. The aforementioned selectivity, however, is violated due to the HF mixing of states at the intermediate i-step, which results in uncontrolled population leak (or information losses in applied problems) into other undesirable HF sublevels.
@@
Further analysis presented in subsection IV.2 and Appendix C demonstrates that introduction of a third, “blocking” laser (control laser) enables active switching off of some off-diagonal (mixing) HF terms presented in Tables 23, and therefore allows tuning between the $|\Delta F|\leq 1$ and $|\Delta F|\equiv 0$ regimes of two-photon excitation. As our numerical calculations for Na and Rb atoms have demonstrated, the deviation from selective population of HF components achieved in this way can be less than $0.001\%$ (see Fig.13). The control laser method introduced here thus opens a “window of opportunity” for manipulating the intra-atomic interactions, in our case eliminating the HF effects even in Rb atoms with a considerably strong HF coupling.
@@ B.1 Rabi frequencies and atomic parameters
where $\Phi=I+J+1$ and $I=3/2$ is the nuclear spin for both Na and Rb atoms. The symbol $(nlJ\parallel D\parallel n^{\prime}l^{\prime}J^{\prime})$ gives the reduced matrix element for the involved fine transition $\{lJ\to l^{\prime}J^{\prime}\}$ (the electron spin is $s=1/2$). The above expressions (26-27) unequally define the dipole interaction operator $\hat{V}^{(P,S)}$ under RWA for the given values of laser field amplitudes $E_{P,S}$, or equally, for the characteristic frequencies $\Omega_{P,S}$ (26).
@@ Appendix C Selectivity assessment under the presence of a control laser
If we take into account now that the two-photon excitation of the desired Rydberg HF component ${\left|f\right.\rangle}_{2}$ goes via the state ${\left|D\right.\rangle}\!\simeq\!{\left|i\right.\rangle}_{2}$ (see Fig.14), the corresponding effective Rabi frequency can be roughly approximated as

### 24 Dec 2023 v2

@@ I Introduction
Although a variety of coherent techniques such as the stimulated Raman adiabatic passage technique (STIRAP) Vitanov et al. (2017); Shore (2017) have enabled nearly $100\%$ efficient excitation of the selected target quantum state Garcia-Fernandez et al. (2005), for alkali-metal atoms the perfect level addressing problem remains unsolved due to existence of Hyperfine (HF) splitting of the atoms’ energy levels. HF structure of highly excited levels with large principal quantum numbers $n$ can not be resolved due to the small energy separation ($\sim n^{-3}$) between sublevels. Thus no separate HF sublevel can be populated by any tuning of the exciting laser. Our research presented here (see Sections II and IV.2) demonstrates that selective excitation can still be achieved for a few two-photon excitation schemes of type $3(5)s_{1/2}-3(5)p_{3/2,1/2}-ns_{1/2}$ in ${}^{23}$a and ${}^{87}$b atoms (see Fig. 1). Throughout this paper, the principal quantum numbers and the HF energy level splitting values placed in parentheses refer to rubidium atoms. In these excitation schemes, the total angular momentum ${F}^{\prime\prime}$ of a well-resolved ground $3(5)s_{1/2}$ state has to remain the same for the final $ns_{1/2}$ state. As long as the P-laser couples all HF components $F^{\prime}$ of a $3(5)p_{3/2,1/2}$ state in the first excitation step ($\Delta F_{PS}=0,\pm 1$ according to the one-photon selection rule), while in the second step the S-laser couples both HF components $F=1,2$ of $ns_{1/2}$ state ($\Delta F_{SP}=0,\pm 1$), one of which turns out to be unpopulated, we can speak of a specific two-photon selection rule $\Delta F_{SPS}\equiv 0$ observed instead of the expected $\Delta F_{SPS}=0,\pm 1$.
@@ II HF blocks architecture of two-step excitation schemes
Dynamics of atomic systems under the influence of periodic or polychromatic control fields can be accurately described using various modifications of the Floquet technique Chu and Telnov (2004); Efimov et al. (2014); Eckardt and Anisimovas (2015). A useful tool for theoretical analysis in the case of nearly–resonant monochromatic laser fields $\textbf{E}\cos(\omega_{L}t)$ is the rotating wave approximation (RWA) Shore (2009). In this framework, the problems of light-matter interactions are reduced to solving the quasistationary optical Bloch equations for evolution of the atomic density matrix, with characteristic timescales being determined by the duration of laser pulses, i.e. the temporal behavior of the electrical field amplitudes $\textbf{E}(t)$. A key component of the theoretical analysis is the concept of adiabatic (“laser-dressed”) atomic state basis, which often allows to develop quantitative description of the phenomena at hand for relatively simple systems and to predict possible evolution scenarios in the case of multidimensional configurations Shore (2017); Rangelov et al. (2006). The latter have been considered as a promising physical objects for quantum processing on the so-called qudits (multidimensional extension of qubits), permitting to design effective algorithms for the implementation of universal gates via adiabatic passage processes Rousseaux et al. (2013); Cinins et al. (2022). In this section we develop an approach that enables identifying dressed states in multidimensional two-step ladder excitation schemes, in the specific case of alkali metal atoms interacting with linearly polarized laser light.
@@ II.2 Dipole matrix elements
The relation (5) indicates that both laser fields couple only the levels with the same quantum numbers $m_{J},m_{I}$, and the corresponding linkage diagram shown in Fig. 3 is reduced to a set of separate, non-interacting (independent) three-level ladders
@@ II.3 Building the Morris-Shore basis vectors
The arbitrarily chosen basis vectors $|d_{M\xi}\rangle$ ($\xi\!=\!0,...,\kappa$) in the manifold $\Lambda_{nd_{5/2}}^{DS}$ are in no way coupled with either the ground or intermediate states (see Fig.2), i.e. are “Dark”.
@@ III.1 Formation of multiplet structure in AT signals
Figure 4: Energies $\varepsilon$ of adiabatic (dressed) states in Na vs the S-laser reduced Rabi frequency $\Omega_{S}$ (26) for the $3s_{1/2}(F^{\prime\prime},M)\!\to\!3p_{3/2}(F^{\prime},M)\!\to\!4d_{5/2}(F,M)$ excitation Zeeman sequences ($M\!=\!0,1,2$) in the case of pump laser detuning $\Delta_{S}\!=\!-30\,\text{MHz}$. The dashed vertical lines correspond to $\Omega_{S}$-values presented in Fig. 5 frames (a)-(d). The square brackets indicate the states that transfer into dark states. The zero-energy positions related to the energy of the $3p_{3/2}(F^{\prime}\!=\!2)$-state. The curves labeled with quantum numbers $F^{\prime}$ or $F$ refer to the dressed states starting to evolve at $\Omega_{S}=0$ from the corresponding HF basis vectors ${\left|\gamma,F,M\right.\rangle}$ ($\gamma\!=\!i,f$).
@@
Strong coupling implies that both $\Omega_{1/2,3/2}^{(S)}$ exceed the bare energy separations $\Delta{\widetilde{\varepsilon}}_{\varsigma}$ (14) for all coupling pairs, so that at large reduced frequency $\Omega_{S}$ (see its definition in Appendix B, Eq. (26)) relation (14) reduces to the simple linear form:
@@
Figure 6: (Color online) The same as in Fig.4 in the case of $3s_{1/2}(F^{\prime\prime}=2)\to 3p_{3/2}\to 5s_{1/2}$ excitation sequence in Na for different $S$-laser reduced Rabi frequencies $\Omega_{S}$. The corresponding RWA linkage diagrams are shown in frames (a1), (b1), (c1). Simulations are performed using parameters $\Omega_{P}=0.87\,\text{MHz}$, $\Delta_{S}=-115\,\text{MHz}$.
@@ III.2 Numerical simulation
Figure 7: (Color online) The same as in Fig.5 in the case of $3s_{1/2}(F^{\prime\prime}=2)\to 3p_{3/2}\to 5s_{1/2}$ excitation sequence for different $S$-laser reduced Rabi frequencies $\Omega_{S}$. Simulations are performed for Na atoms using parameters $\Omega_{P}=0.87\,\text{MHz}$, $\Delta_{S}=-115\,\text{MHz}$. The expected AT peaks (the vertical lines) are consistent with the data in Fig.6.
@@ IV.1 Spectral composition and selection rules
With regard to the spectra of Figs. 5, 7, one cleanly sees both signs of spectra pauperisation, namely: (a) multiplet structure and (b) lack of all AT multiplets $1_{\pm}$ related to the excitation scheme $3s_{1/2}(F^{\prime\prime}\!=\!1)\rightarrow 3p_{3/2}$. Bright multiplets $2_{\pm}$ and $1_{\pm}$ turn out to complement each other, i.e. their visualization occurs during independent probing from different initial HF components $F^{\prime\prime}\!=\!2,1$ of the ground state. The data in Table 2(a) do show that HF interaction poorly couples all ${\left|i\right.\rangle}_{1}$ and ${\left|i\right.\rangle}_{2}$ BS: the corresponding $\Omega_{HF}$ rates are $\Omega_{HF}\!=\!0$ for $M\!=\!0,2$ while for two BS with $M\!=\!1$ the sharing parameter (17) $R\!=\!(7.44/65.6)^{2}\!=\!0.0129$. Practically the same thing happens in the case of Rb atoms, as can be seen from the corresponding data in Table 3(a): for both $M\!=\!0,2$ the values of $\Omega_{HF}\!=\!0$ and the the sharing parameter (17) $R\!=\!(34.0/299.7)^{2}\!=\!0.0128$ for two BS with $M\!=\!1$.
@@ V Conclusions
An interesting feature of the excitation ladders $ks_{1/2}\to kp_{3/2,1/2}\to ns_{1/2}$ manifests as preservation of the HF quantum number $F$ in transitions between the initial ground $ks_{1/2}$ and the final $ns_{1/2}$ states, resulting in a specific selection rule $\Delta F\!=\!F^{\prime\prime}\!-\!F=0$ for the two-photon transitions $ks_{1/2}(F^{\prime\prime})\to ns_{1/2}(F)$. The reason for this phenomenon is related to the physics of the corresponding excitation processes proceeding through a series of partial two-photon paths with different intermediate HF sublevels (see Fig. 6 (b1)). The destructive interference between probability amplitudes forbids the two-photon transition when $F_{f}\neq F_{g}$. This observation opens practically important perspectives for artificially introducing custom selection rules in many-photon interactions both to manipulate quantum states and to achieve individual addressing of HF components of atomic and molecular energy levels.
@@
In the absence of HF interaction, the control radiation of P- and S-lasers allows ideal ($100\%$) selective excitation of single three-level ladder blocks (10), i.e. excitation of selected bright $\eta$-states with a fixed $\eta$-index ($\eta=F^{\prime\prime}M$). With them, for example, using the STIRAP technique, it is possible to carry out some basic quantum operations Vitanov et al. (2017) or, on the basis of these three $\eta$-states, to create independent cells for storing optical information by forming proper $\eta$-polaritons Vitanov et al. (2017). Coherent excitation of Rb Rydberg states in a three-level Levine et al. (2019) or four-level Beterov et al. (2023) elementary ladder blocks with $F^{\prime\prime}=2,M=0$ is an important element of a three-qubit Toffoli gate implementation Levine et al. (2019). The data presented in Table 3 provides useful information for determining the optimal parameters (Rabi frequencies and laser detunings) in this type of experiments Levine et al. (2019); Beterov et al. (2023) which use high-contrast Rydberg pulses.
@@
The aforementioned selectivity, however, is violated due to the HF mixing of states at the intermediate i-step, which results in uncontrolled population leak (or information losses in applied problems) into other undesirable HF sublevels. Further analysis presented in subsection IV.2 and Appendix C demonstrates that introduction of a third, “blocking” laser (control laser) enables active switching off of some off-diagonal (mixing) HF terms presented in Tables 23, and therefore allows tuning between the $|\Delta F|\leq 1$ and $|\Delta F|\equiv 0$ regimes of two-photon excitation. As our numerical calculations for Na and Rb atoms have demonstrated, the deviation from selective population of HF components achieved in this way can be less than $0.001\%$ (see Fig.13). The control laser method introduced here thus opens a “window of opportunity” for manipulating the intra-atomic interactions, in our case eliminating the HF effects even in Rb atoms with a considerably strong HF coupling.
@@ B.1 Rabi frequencies and atomic parameters
where $\Phi=I+J+1$ and $I=3/2$ is the nuclear spin for both Na and Rb atoms. The symbol $(nlJ\parallel D\parallel n^{\prime}l^{\prime}J^{\prime})$ gives the reduced matrix element for the involved fine transition $\{lJ\to l^{\prime}J^{\prime}\}$ (the electron spin is $s=1/2$). The above expressions (26-27) unequally define the dipole interaction operator $\hat{V}^{(P,S)}$ under RWA for the given values of laser field amplitudes $E_{P,S}$, or equally, for the reduced Rabi frequencies $\Omega_{P,S}$ (26).
@@ Appendix C Selectivity assessment under the presence of a control laser
If we take into account now that the two-photon excitation ${\left|g\right.\rangle}_{2}\!\to\!{\left|f\right.\rangle}_{2}$ of the desired Rydberg HF component ${\left|f\right.\rangle}_{2}$ at a two-photon resonance goes via the state ${\left|D\right.\rangle}\!\simeq\!{\left|i\right.\rangle}_{2}$ (see Fig.14 at the same energies of quantum states ${\left|f\right.\rangle}_{2}$ and ${\left|g\right.\rangle}_{2}$), the corresponding effective Rabi frequency can be roughly approximated as